SapphireVoid:

∠1 = 20x + 21 and ∠2 = 30x − 29

snowflake0531:

IS there a graph that goes with it?

SapphireVoid:

no

SapphireVoid:

but i can us a screen shot

SapphireVoid:

of the angles

SapphireVoid:

1 attachment
snowflake0531:

Yes Part A is correct 20x+21 = 30x -29 You want to first isolate the x value, so I suggest adding 29 to both sides first o-O 20x+21 +29 = 30x-29+29 what does that simplify to?

SapphireVoid:

ok

SapphireVoid:

side A 20x + 50 side B 30x

snowflake0531:

20x+50=30x Now do 20x-20x+50 = 30x-20x

SapphireVoid:

ok

SapphireVoid:

side A 50 side B 10x

SapphireVoid:

oops didnt mean to put the x

snowflake0531:

yes 50=10x So what is 50/10

SapphireVoid:

5

snowflake0531:

So now we established that x is 5 we know that hte angles are equal so just ifnd one of them 20x+21 20(5)+21

SapphireVoid:

ok

SapphireVoid:

121

snowflake0531:

yep so there's your answer for both

SapphireVoid:

ok thannk you

snowflake0531:

yw~